OpenStudy (anonymous):

4 and 8/10 divided by 4/5

OpenStudy (anonymous):

Change 4 and 8/10 into an improper fraction: ((4*10)+8)/10 = 48/10 When you divide fractions, you take the reciprocal (flip) the divisor (the number you're dividing by) and multiply it by your dividend. In your case, 48/10 / 4/5 = 48/10 * 5/4 = (48*5)/(10*4) = 240/40 = 6